# Design and Performance Simulation of the Electromagnetic Calorimeter at EicC

### v1

@@ 1 Introduction
The Electron-Ion Collider in China (EicC) [bib:eiccWhitePaper], proposed as a new facility at the High Intensity Heavy-ion Accelerator Facility (HIAF) [YANG2013263] in Huizhou, aims to advance our understanding of quantum chromodynamics (QCD). Through high-energy electron-ion collisions, it will investigate the internal structure of nucleons, the dynamics of quarks and gluons, and the fundamental nature of the strong force.
@@
The Electromagnetic Calorimeter (ECAL) of EicC is responsible for detecting electrons and photons within a pseudorapidity range of , corresponding to approximately 99% of the total solid angle. To achieve the required performance across this coverage, the ECAL is divided into three segments: the electron-endcap (e-Endcap), the central barrel, and the ion-Endcap. Each segment adopts a calorimeter technology optimized for its specific physics and geometrical constraints.
@@ 3.3 Electron and Pion Identification
One of the key performance requirements of the EicC ECAL is to provide additional hadron rejection capability, particularly for distinguishing pions from electrons, while simultaneously maintaining high efficiency for electron detection. Electrons typically deposit nearly all of their energy in the calorimeter via the development of an electromagnetic shower. In contrast, high-energy hadrons primarily behave as Minimum Ionizing Particles (MIPs), depositing only a fraction of their kinematic energy in the calorimeter. This difference can be exploited using the ratio, defined as the energy deposited in the calorimeter divided by the momentum measured by the tracking detector. For this simulation, the momentum is taken from the Monte Carlo truth information.
@@ 5 Conclusions
A comprehensive design and optimization study of the Electromagnetic Calorimeter for the EicC has been presented. The ECAL integrates two complementary technologies to meet distinct physics and geometric requirements: a pCsI crystal calorimeter for the e-Endcap and a Shashlik sampling calorimeter for the barrel and ion-Endcap regions.
Simulation results demonstrate that both systems achieve the target energy and position resolutions essential for precise electron and photon measurements. The pCsI modules provide excellent energy resolution and timing performance, benefiting from their high light yield and ultraviolet response. The Shashlik modules, optimized for radiation length, sampling structure, and light collection uniformity, deliver reliable performance and efficient separation in a cost-effective design.
Together, the two calorimeter technologies provide comprehensive coverage across the detector acceptance, supporting precise energy measurement, particle identification, and reconstruction. Future work will focus on prototype construction, experimental validation, and integration into the EicC detector system.

### v2

@@ 1 Introduction
The Electron-Ion Collider in China (EicC) [bib:eiccWhitePaper], proposed as a new facility at the High Intensity Heavy-ion Accelerator Facility (HIAF) [YANG2013263] in Huizhou, aims to advance our understanding of QCD. Through high-energy electron-ion collisions, it will investigate the internal structure of nucleons, the dynamics of quarks and gluons, and the fundamental nature of the strong force.
@@
The ECAL of EicC is responsible for detecting electrons and photons within a pseudorapidity range of , corresponding to approximately 99% of the total solid angle. To achieve the required performance across this coverage, the ECAL is divided into three segments: the electron-endcap (e-Endcap), the central barrel, and the ion-Endcap. Each segment adopts a calorimeter technology optimized for its specific physics and geometrical constraints.
@@ 3.3 Electron and Pion Identification
One of the key performance requirements of the EicC ECAL is to provide additional hadron rejection capability, particularly for distinguishing pions from electrons, while simultaneously maintaining high efficiency for electron detection. Electrons typically deposit nearly all of their energy in the calorimeter via the development of an electromagnetic shower. In contrast, high-energy hadrons primarily behave as Minimum Ionizing Particles (MIPs), depositing only a fraction of their kinematic energy in the calorimeter. This difference can be exploited using the ratio, defined as the energy deposited in the calorimeter divided by the momentum measured by the tracking detector. For this simulation, the momentum is taken assuming ideal tracking resolution.
@@ 5 Conclusions
A comprehensive design and optimization study of the ECAL for the EicC has been presented, combining two complementary technologies to meet the distinct physics and geometric requirements of the detector: a pCsI crystal calorimeter in the e-Endcap and Shashlik style sampling calorimeters in the barrel and ion-Endcap regions.
Detailed Geant4 simulations, incorporating realistic detector responses, show that the proposed configuration meets the performance targets required by the EicC physics program. The pCsI modules offer excellent energy resolution of about , superior position resolution, and strong timing performance, benefiting from their high light yield. The Shashlik modules, optimized in sampling structure, radiation length, and light collection uniformity, provide energy and position resolutions of approximately and 5 mm at 1 GeV, respectively, along with effective electron–pion separation.
These results indicate that the current ECAL design can reliably support precision electron and photon measurements across the full acceptance of the EicC detector. Future work will focus on detailed digitization modeling, prototype construction and validation, and advancing the ECAL toward full engineering design and construction.